May 25th UT LBC Tech Night 2

Observers: Pedani, Kuhn OSA: Gonzalez Huerta

02:05 Starting taking 25 Biases. Outside T_air=8.4, T_mir=7.9 Scattered Clouds

05:27 Dome still shut - even some rain drops. We decided to try 'mock' observing, but find UI hung and IIFGUI says mount and rotator "not ready". Observer error... did not realize the browse box from the UI was open already. Cannot even claim it was the altitude. OB of PC1643+4631A through dome completed OK.

09:00 Another set of 10 biases. For some reason both blue and red PAs are at 320deg, although OB commands PA to 0 deg. IIFGUI says mount & rotator "not ready", mount "not ready" and preset "canceled". Telescope was parked at zenith, that's why. Slewing and repeating set of biases. Now tracking, PA is 0.

09:30 Clouds continue - we are closing.

-- MarcoPedani - 25 May 2009
I Attachment Action Size Date Who Comment
daq.20090525.gifgif daq.20090525.gif manage 34 K 17 Jul 2009 - 05:02 JohnHill Plot of telescope and mirror temperatures
Topic revision: r5 - 17 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback